# Neutrino Constraints on Black Hole Formation in M31

### v1

@@ 1 Introduction
Recently, the disappearance of a massive star, identified as M31-2014-DS1, in the Andromeda galaxy (M31) was reported without a typical luminous supernova explosion (De et al., 2024). Such phenomena, known as “failed supernovae,” involve the black hole formation via gravitational collapse without a significant optical outburst. Theoretical studies have long predicted that certain massive stars could end their evolution through a quiet collapse directly into black holes, particularly stars characterized by highly compact cores or masses exceeding specific thresholds (O’Connor & Ott, 2011; Sukhbold et al., 2016). Systematic numerical simulations by O’Connor & Ott (2011) demonstrated a strong correlation between core compactness and the likelihood of failed explosions, emphasizing that stars exceeding a critical value of the compactness parameter are likely to collapse silently into black holes rather than undergoing successful explosions.
@@
From 2008 to 2018, the Super-Kamiokande (SK) neutrino observatory conducted nearly a decade of continuous neutrino monitoring, actively searching for astrophysical neutrino bursts associated with gravitational collapse events (Mori et al., 2022). Although no significant neutrino signals coinciding with the timing of M31-2014-DS1 were detected, these null detections still provide important constraints on theoretical models of neutrino emissions and core-collapse mechanisms. Present neutrino detectors, including SK, primarily possess sensitivity limited to events within or near our galaxy, typically within hundreds of kpc. Therefore, while the non-detection from M31 is consistent with current detector capabilities, SK could have had a 20%-40% chance of detection at the distance of M31 assuming the neutrino emission model of Nakazato et al. (2010). This emphasizes the necessity for future detectors, such as Hyper-Kamiokande, JUNO, and DUNE, designed to improve detection sensitivity to extragalactic core-collapse neutrinos. Upcoming detectors may feasibly detect neutrino emissions from events like M31-2014-DS1, significantly advancing our ability to study stellar collapse physics and the properties of black hole formation.
@@ 2 Neutrino Signatures Predicted by Black Hole Formation Simulations
References [^1] In this study, we only use simulation results that include comprehensive neutrino data. Progenitor mass EOS Gravity $t_{\rm BH}$222postbounce time $E_{\nu_{e}}$ $E_{\bar{\nu}_{e}}$ $E_{\nu_{X}}$ $\varepsilon_{\nu_{e}}\right>$ $\varepsilon_{\bar{\nu}_{e}}\right>$ $\varepsilon_{\nu_{X}}\right>$
@@
Sumiyoshi et al. (2007) 40 (WW95333Woosley & Weaver (1995)) LS180 fGR444fully GR 0.56 0.554 0.467 0.228 16.3 19.5 21.5
@@
Sumiyoshi et al. (2008) 40 (H95555Hashimoto (1995)) LS180 fGR 0.36 0.334 0.271 0.160 13.5 16.8 21.9
Sumiyoshi et al. (2008) 50 (TUN07666Tominaga et al. (2007)) Shen-TM1 fGR 1.51 1.35 1.27 0.526 20.0 23.1 24.2
@@
Fischer et al. (2009) [^7] See also Liebendörfer et al. (2004) 40 (WW95) LS180 fGR 0.435 0.507 0.376 0.231 14.1 14.6 19.7
@@
Walk et al. (2020) 40 (WH07101010Woosley & Heger (2007)) LS220 eGR 0.57 0.572 0.539 0.375 16.2 18.8 20.2
Nakazato et al. (2021) 30 (N13111111Nakazato et al. (2013)) LS220 fGR 0.342 0.403 0.287 0.211 12.5 16.4 22.3
@@
Choi et al. (2025) [^12] See also Burrows et al. (2024) 12.25 (S16&18131313Sukhbold et al. (2016, 2018)) SFHo eGR $>$2.09 0.563 0.511 0.297 13.9 16.3 15.5
Choi et al. (2025) 14 (S16&18) SFHo eGR $>$2.82 0.768 0.711 0.393 15.0 17.3 15.9
@@
unpublishded [^14] The same code as Nakazato et al. (2010) 40 (WW95) Togashi fGR 0.927 0.824 0.705 0.471 18.1 20.7 25.7
@@
Table 1 gives a summary of simulation results for core-collapse supernovae resulting in black hole formation from various studies. Columns show the references, progenitor mass, the employed equation of state (EOS), treatment of gravity (fGR for fully general relativistic and eGR for effectively general relativistic; the latter employs Newtonian hydrodynamics combined with an approximate gravitational potential incorporating general relativistic effects), black hole formation time from core bounce ($t_{\mathrm{BH}}$), emitted neutrino energies ($E_{\nu_{e}}$, $E_{\bar{\nu}_{e}}$, $E_{\nu_{X}}$), and corresponding mean neutrino energies ($\langle\varepsilon_{\nu_{e}}\rangle$, $\langle\varepsilon_{\bar{\nu}e}\rangle$, $\langle\varepsilon_{\nu_{X}}\rangle$). Simulation outcomes reveal a systematic dependence of the neutrino signal on the stiffness of the EOS. Stiffer EOSs sustain longer emission phases, yielding larger integrated neutrino energies and correspondingly higher expected event counts. In contrast, softer EOSs terminate the emission sooner, producing smaller total energies and fewer events (Sumiyoshi et al., 2006). Although progenitor mass modulates these trends – more massive stars generally enhance the overall neutrino output – the distinction between stiff and soft EOSs remains the primary factor governing both the duration and the amplitude of the predicted neutrino signal.
@@
The EOS significantly affects neutrino emission by determining the thermodynamic and nuclear properties of dense matter during collapse. Several EOS models have been employed in the simulations, which are summarized here. The Lattimer-Swesty EOS (Lattimer & Swesty, 1991), based on a Skyrme-type nuclear interaction and a compressible liquid-drop model, has been widely used in earlier simulations. This EOS includes variations characterized by different nuclear incompressibility parameters ($K$); among the simulations presented in this work, only the $K=180$ MeV (LS180) and $K=220$ MeV (LS220) variants are used. The Shen EOS family (Shen et al., 1998, 2020), constructed within relativistic mean-field theory, is represented here by two parameter sets: the original one (Shen-TM1) and the updated one (Shen-TM1e). Shen-TM1e adopts a reduced symmetry-energy slope compared with Shen-TM1, resulting in a slightly softer pressure at supra-nuclear densities and hence altered neutrino luminosities and spectra. The Togashi EOS (Togashi et al., 2017) employs advanced nuclear many-body calculations, offering detailed composition and structure predictions at high densities. Lastly, the SFHo EOS (Steiner et al., 2013), derived from relativistic mean-field models constrained by experimental nuclear data, is frequently utilised owing to its realistic treatment of nuclear-matter properties at densities relevant to PNS formation and neutrino emission. Since temperatures during black hole formation are very high, it is essential to consider finite-temperature effects rather than zero-temperature properties when comparing EOS models. In this analysis, we adopt an entropy value of $4\,k_{\mathrm{B}}$/baryon as a typical reference condition to evaluate each EOS consistently under relevant collapse conditions. Under these finite-temperature conditions (entropy of 4 $k_{\mathrm{B}}$/baryon), the maximum gravitational masses of PNSs, $M_{s=4}^{\rm max}$, predicted by each EOS are approximately: LS180 (2.0 $M_{\odot}$), LS220 (2.1 $M_{\odot}$), SFHo (2.3 $M_{\odot}$), Togashi (2.4 $M_{\odot}$), Shen-TM1e (2.5 $M_{\odot}$) and Shen-TM1 (2.6 $M_{\odot}$) (see, e.g., Steiner et al., 2013; da Silva Schneider et al., 2020). Because the maximum mass increases with the stiffness of the EOS, a stiffer EOS not only supports more massive PNSs but also tends to yield higher neutrino luminosities during collapse, thereby boosting the number of neutrino events expected in detectors. Comparing simulation results across these EOS models helps constrain uncertainties in nuclear physics that impact neutrino emission modeling from stellar collapses.
@@ 3 Expected Neutrino Events and Constraints from Super-Kamiokande
Figure 1: Total emitted electron antineutrino energy $E_{\bar{\nu}_{e}}$ versus mean energy $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$ predicted by black-hole-forming core-collapse simulations. The shaded bands indicate regions that would have produced $\geq 2$ correlated events in SK with probabilities of 50 %, 68 %, 95 %, and 99 % (light to dark blue), assuming Poisson statistics for a source at the distance of M31. Filled circles show individual simulation results for six nuclear equations of state (EOS): LS180, LS220, SFHo, Togashi, Shen-TM1e, and Shen-TM1. Both grey tone and symbol size encode the maximum gravitational mass of a non-rotating protoneutron star supported by each EOS: lighter symbols correspond to softer EOSs with the maximum gravitational mass of protoneutron star with $s=4k_{B}/\text{baryon}$, $M^{\max}_{s=4}=2.0\text{--}2.1\,M_{\odot}$ (LS180, LS220), whereas darker symbols mark stiffer EOSs with $M^{\max}_{s=4}=2.5\text{--}2.6\,M_{\odot}$ (Shen-TM1e, Shen-TM1). Models that lie outside the shaded regions, where fewer than two events would be expected, remain compatible with the SK non-detection, illustrating the strong dependence of the predicted neutrino signal on EOS stiffness and the associated maximum neutron-star mass.
@@
[…]
@@
where $M_{\rm det}$ is the detector mass (22.5 kton, corresponding to the fiducial volume of SK), $m$ is the nucleon mass, $D$ is the distance between the supernova and Earth (770 kpc, corresponding to the distance to M31), and $\sigma$ is the inverse beta decay cross-section ($p+\bar{\nu}_{e}\to n+e^{+}$). [^15] In this calculation, we focus exclusively on inverse beta decay (IBD) interactions. Although the observational limits derived by Mori et al. (2022) were obtained considering all neutrino interactions within the relevant energy window, contributions from channels other than IBD are relatively minor. Thus, limiting our estimate to IBD provides a reasonable approximation. For the cross-section, we adopt $\sigma(\varepsilon_{\nu})=\sigma_{0}(\varepsilon_{\nu}/{\rm MeV})^{2}$ with $\sigma_{0}=9.4\times 10^{-44}\,{\rm cm}^{2}$ (Raffelt, 1996) and assume a thermal neutrino spectrum (see Section 4 in Suwa et al. 2021 for more details). [^16] In this calculation, a purely thermal neutrino spectrum is assumed. However, a nonthermal neutrino component might also arise due to compressional flows onto the newly formed black hole or turbulent motion (Suwa, 2013; Suwa et al., 2019; Nagakura & Hotokezaka, 2021), potentially increasing the detectability significantly. Precise estimates should consider such effects, as well as detailed detector responses and efficiencies. The factor $2/18$ accounts for the number of hydrogen atoms per water molecule, making $2M_{\rm det}/18m$ the total number of hydrogen nuclei within the detector volume.
@@
Using ten years of SK‐IV data (2008–2018), Mori et al. (2022) searched for neutrino bursts from distant core-collapse supernovae with a statistical clustering algorithm. To suppress accidental coincidences, the analysis required either (i) at least two candidate events within a 0.5 s or 2 s window, or (ii) at least four events within a 10 s window. Despite a background rejection rate exceeding $99.3\,\%$ and a detection efficiency of roughly $74\,\%$, no neutrino cluster passed these criteria. The energy threshold was set at 5.5 MeV, a value low enough that it does not limit the detection of the higher-energy neutrinos expected from failed supernovae.
@@
Figure 1 shows the relationship between the total emitted energy by anti-electron type neutrinos $E_{\bar{\nu}_{e}}$ and the mean neutrino energy $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$. The shaded regions represent excluded parameter space based on the non-detection of correlated neutrino events in SK, assuming Poisson statistics. Each contour corresponds to a confidence level of 50%, 68%, 95%, or 99%, and indicates the set of emission parameters ($E_{\bar{\nu}_{e}}$, $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$) for which the probability of detecting two or more events would have exceeded the given confidence level. For example, an expected number of approximately 1.68 events corresponds to a 50% chance of detecting two or more events, while about 2.35, 4.74, and 6.64 expected events correspond to the 68%, 95%, and 99% confidence levels, respectively.
@@
$F_{\bar{\nu}_{e}}^{\mathrm{obs}}=\begin{cases}\cos^{2}\theta_{12}\,F_{\bar{\nu}_{e}}^{\mathrm{src}}+\sin^{2}\theta_{12}\,F_{\bar{\nu}_{X}}^{\mathrm{src}}&\text{(normal ordering)}\\ F_{\bar{\nu}_{X}}^{\mathrm{src}}&\text{(inverted ordering)}\end{cases}.$ (2)
@@
Different EOS models – LS180 ($M_{s=4}^{\rm max}=2.0,M_{\odot}$), LS220 ($M_{s=4}^{\rm max}=2.1,M_{\odot}$), SFHo ($M_{s=4}^{\rm max}=2.3,M_{\odot}$), Togashi ($M_{s=4}^{\rm max}=2.4,M_{\odot}$), Shen-TM1e ($M_{s=4}^{\rm max}=2.5,M_{\odot}$) and Shen-TM1 ($M_{s=4}^{\rm max}=2.6,M_{\odot}$) – are represented by symbols with varying shades and sizes. Most simulation results cluster around or below the 50% confidence boundary, indicating good consistency with observational constraints, particularly at lower neutrino energies. The Shen-TM1 EOS, associated with the largest maximum PNS mass ($2.6M_{\odot}$), predicts higher total neutrino energies, yet these results also fall within the observational limits. This clearly demonstrates how neutrino emission properties depend on the chosen EOS model.

### v2

@@ 1 Introduction
Recently, the disappearance of a massive star, identified as M31-2014-DS1, in the Andromeda galaxy (M31) was reported without a typical luminous supernova explosion (De et al., 2024). Such phenomena, often referred to as “failed supernovae,” involve black hole formation via gravitational collapse without a significant optical outburst. Theoretical studies have long predicted that certain massive stars could end their evolution through a quiet collapse directly into black holes, particularly stars characterized by highly compact cores or masses exceeding specific thresholds (O’Connor & Ott, 2011; Sukhbold et al., 2016). Systematic numerical simulations by O’Connor & Ott (2011) demonstrated a strong correlation between core compactness and the likelihood of failed explosions, emphasizing that stars exceeding a critical value of the compactness parameter are likely to collapse silently into black holes rather than undergoing successful explosions.
@@
From 2008 to 2018, the Super-Kamiokande (SK) neutrino observatory conducted nearly a decade of continuous neutrino monitoring, actively searching for astrophysical neutrino bursts associated with gravitational collapse events (Mori et al., 2022). Although no significant neutrino signals coinciding with the timing of M31-2014-DS1 were detected, these null detections still provide important constraints on theoretical models of neutrino emissions and core-collapse mechanisms. Present neutrino detectors, including SK, primarily possess sensitivity limited to events within or near our galaxy, typically within hundreds of kpc. Therefore, while the non-detection from M31 is consistent with current detector capabilities, SK could have had a 20%-40% chance of detection at the distance of M31 assuming the neutrino emission model of Nakazato et al. (2010). This emphasizes the necessity for future detectors, such as Hyper-Kamiokande, JUNO, and DUNE, designed to improve detection sensitivity to extragalactic core-collapse neutrinos. Upcoming detectors may feasibly detect neutrino emissions from events such as the failed supernova candidate M31-2014-DS1, significantly advancing our ability to study stellar collapse physics and the properties of black hole formation.
@@ 2 Neutrino Signatures Predicted by Black Hole Formation Simulations
References [^1] In this study, we only use simulation results that include comprehensive neutrino data. Progenitor mass EOS Gravity $t_{\rm BH}$ [^2] Postbounce time. $E_{\nu_{e}}$ $E_{\bar{\nu}_{e}}$ $E_{\nu_{X}}$ $\left<\varepsilon_{\nu_{e}}\right$> $\left<\varepsilon_{\bar{\nu}_{e}}\right$> $\left<\varepsilon_{\nu_{X}}\right$>
@@
Sumiyoshi et al. (2007) 40 (WW95333Woosley & Weaver (1995).) LS180 fGR [^4] Fully GR. 0.56 0.554 0.467 0.228 16.3 19.5 21.5
@@
Sumiyoshi et al. (2008) 40 (H95555Hashimoto (1995).) LS180 fGR 0.36 0.334 0.271 0.160 13.5 16.8 21.9
Sumiyoshi et al. (2008) 50 (TUN07666Tominaga et al. (2007).) Shen-TM1 fGR 1.51 1.35 1.27 0.526 20.0 23.1 24.2
@@
Fischer et al. (2009) [^7] See also Liebendörfer et al. (2004). 40 (WW95) LS180 fGR 0.435 0.507 0.376 0.231 14.1 14.6 19.7
@@
Walk et al. (2020) 40 (WH07101010Woosley & Heger (2007).) LS220 eGR 0.57 0.572 0.539 0.375 16.2 18.8 20.2
Nakazato et al. (2021) 30 (N13111111Nakazato et al. (2013).) LS220 fGR 0.342 0.403 0.287 0.211 12.5 16.4 22.3
@@
Largani et al. (2024) 40 (WHW02) DD2F RDF-1.7 fGR 1.03 0.368 0.384 0.132 13.3 15.6 14.2
Choi et al. (2025) [^12] See also Burrows et al. (2024), which present three-dimensional simulations. 12.25 (S16&18131313Sukhbold et al. (2016, 2018).) SFHo eGR >>2.09 0.563 0.511 0.297 13.9 16.3 15.5
Choi et al. (2025) 14 (S16&18) SFHo eGR >>2.82 0.768 0.711 0.393 15.0 17.3 15.9
@@
unpublished [^14] The same code as Nakazato et al. (2010). 40 (WW95) Togashi fGR 0.927 0.824 0.705 0.471 18.1 20.7 25.7
@@
[…]
@@
The EOS significantly affects neutrino emission by determining the thermodynamic and nuclear properties of dense matter during collapse. Several EOS models have been employed in the simulations, which are summarized here. The Lattimer-Swesty EOS (Lattimer & Swesty, 1991), based on a Skyrme-type nuclear interaction and a compressible liquid-drop model, has been widely used in earlier simulations. This EOS includes variations characterized by different nuclear incompressibility parameters ($K$); among the simulations presented in this work, only the $K=180$ MeV (LS180) and $K=220$ MeV (LS220) variants are used. The Shen EOS family (Shen et al., 1998, 2020), constructed within relativistic mean-field theory, is represented here by two parameter sets: the original one (Shen-TM1) and the updated one (Shen-TM1e). Shen-TM1e adopts a reduced symmetry-energy slope compared with Shen-TM1, resulting in a slightly softer pressure at supra-nuclear densities and hence altered neutrino luminosities and spectra. The Togashi EOS (Togashi et al., 2017) employs advanced nuclear many-body calculations, offering detailed composition and structure predictions at high densities. The DD2F RDF EOS (Bastian, 2021) is a hadron-quark matter hybrid EOS with a first-order phase transition, and we employ the DD2F RDF-1.7 variant, where 1.7 denotes a specific parameter set in this EOS series characterizing the strength of quark-matter interactions. Lastly, the SFHo EOS (Steiner et al., 2013), derived from relativistic mean-field models constrained by experimental nuclear data, is frequently utilised owing to its realistic treatment of nuclear-matter properties at densities relevant to PNS formation and neutrino emission. Since temperatures during black hole formation are very high, it is essential to consider finite-temperature effects rather than zero-temperature properties when comparing EOS models. In this analysis, we adopt an entropy value of $4\,k_{\mathrm{B}}$/baryon as a typical reference condition to evaluate each EOS consistently under relevant collapse conditions. The corresponding maximum gravitational masses of PNSs, $M_{s=4}^{\rm max}$, predicted by each EOS are approximately: LS180 (2.0 $M_{\odot}$), LS220 (2.1 $M_{\odot}$), SFHo (2.3 $M_{\odot}$), Togashi (2.4 $M_{\odot}$), Shen-TM1e (2.5 $M_{\odot}$) and Shen-TM1 (2.6 $M_{\odot}$) (see, e.g., Steiner et al., 2013; da Silva Schneider et al., 2020). For reference, the DD2F RDF-1.7 EOS yields a similar maximum mass of about 2.1 $M_{\odot}$, where the value at $3\,k_{\mathrm{B}}$/baryon is adopted based on the value reported in the literature. These results indicate that EOSs predicting larger maximum masses also yield longer-lived PNSs. Consequently, such EOSs tend to produce longer neutrino-emission phases and higher total neutrino event counts, whereas those with smaller maximum masses result in earlier emission termination and smaller total counts. The instantaneous luminosity, however, does not necessarily increase with maximum mass, as shown, for example, by Sumiyoshi et al. (2006). Comparing simulation results across these EOS models helps constrain uncertainties in nuclear physics that affect neutrino emission modeling from stellar collapses.
@@ 3 Expected Neutrino Events and Constraints from Super-Kamiokande
Figure 1: Total emitted electron antineutrino energy $E_{\bar{\nu}_{e}}$ versus mean energy $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$ predicted by black-hole-forming core-collapse simulations. The shaded bands indicate regions that would have produced $\geq 2$ correlated events in SK with probabilities of 50 %, 68 %, 95 %, and 99 % (light to dark blue), assuming Poisson statistics for a source at the distance of M31. Filled circles show individual simulation results for seven nuclear equations of state (EOS): LS180, LS220, DD2F RDF-1.7, SFHo, Togashi, Shen-TM1e, and Shen-TM1. Both grey tone and symbol size encode the maximum gravitational mass of a non-rotating protoneutron star supported by each EOS: lighter symbols correspond to softer EOSs with the maximum gravitational mass of protoneutron star with thermal correction, $M^{\max}_{s=4}=2.0\text{--}2.1\,M_{\odot}$ (LS180, LS220), whereas darker symbols mark EOSs with $M^{\max}_{s=4}=2.5\text{--}2.6\,M_{\odot}$ (Shen-TM1e, Shen-TM1). Models that lie outside the shaded regions, where fewer than two events would be expected, remain compatible with the SK non-detection, illustrating the strong dependence of the predicted neutrino signal on EOS stiffness and the associated maximum neutron-star mass.
@@
$N_{\nu}$ $=$ $\frac{2}{18}\frac{M_{\rm det}}{m}\frac{E_{\nu}}{4\pi D^{2}\left<\varepsilon_{\nu}\right>}\left<\sigma\right$> (1)
@@
where $M_{\rm det}$ is the detector mass (22.5 kton, corresponding to the fiducial volume of SK), $m$ is the nucleon mass, $D$ is the distance between the supernova and Earth (770 kpc, corresponding to the distance to M31), and $\sigma$ is the inverse beta decay cross-section ($p+\bar{\nu}_{e}\to n+e^{+}$). [^16] In this calculation, we focus exclusively on inverse beta decay (IBD) interactions. Although the observational limits derived by Mori et al. (2022) were obtained considering all neutrino interactions within the relevant energy window, contributions from channels other than IBD are relatively minor. Thus, limiting our estimate to IBD provides a reasonable approximation. For the cross-section, we adopt $\sigma(\varepsilon_{\nu})=\sigma_{0}(\varepsilon_{\nu}/{\rm MeV})^{2}$ with $\sigma_{0}=9.4\times 10^{-44}\,{\rm cm}^{2}$ (Raffelt, 1996) and assume a thermal neutrino spectrum (see Section 4 in Suwa et al. 2021 for more details). [^17] In this calculation, a purely thermal neutrino spectrum is assumed. However, a nonthermal neutrino component might also arise due to compressional flows onto the newly formed black hole or turbulent motion (Suwa, 2013; Suwa et al., 2019; Nagakura & Hotokezaka, 2021), potentially increasing the detectability significantly. Precise estimates should consider such effects, as well as detailed detector responses and efficiencies. The factor $2/18$ accounts for the mass fraction of free protons in water, making $2M_{\rm det}/18m$ the total number of hydrogen nuclei within the detector volume.
@@
Using ten years of SK‐IV data (2008–2018), Mori et al. (2022) searched for neutrino bursts from distant core-collapse supernovae with a statistical clustering algorithm. To suppress accidental coincidences, the analysis required either (i) at least two candidate events within a 0.5 s or 2 s window, or (ii) at least four events within a 10 s window. Despite a background rejection rate exceeding $99.3\,\%$ and a detection efficiency of roughly $74\,\%$, no neutrino cluster passed these criteria. In this analysis, although the nominal analysis threshold is 5.5 MeV, an effective energy cut of about 15 MeV is applied to suppress background events. Figure 1, described in the following, instead assumes an idealized, background-free detector, representing the ultimate sensitivity rather than the actual SK condition.
@@
Figure 1 shows the relationship between the total emitted energy by electron type antineutrinos $E_{\bar{\nu}_{e}}$ and the mean neutrino energy $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$. The shaded regions represent excluded parameter space based on the non-detection of correlated neutrino events in SK, assuming Poisson statistics. Each contour corresponds to a confidence level of 50%, 68%, 95%, or 99%, and indicates the set of emission parameters ($E_{\bar{\nu}_{e}}$, $\langle\varepsilon_{\bar{\nu}_{e}}\rangle$) for which the probability of detecting two or more events would have exceeded the given confidence level. For example, an expected number of approximately 1.68 events corresponds to a 50% chance of detecting two or more events, while about 2.35, 4.74, and 6.64 expected events correspond to the 68%, 95%, and 99% confidence levels, respectively.
@@
$F_{\bar{\nu}_{e}}^{\mathrm{osc}}=\begin{cases}\cos^{2}\theta_{12}\,F_{\bar{\nu}_{e}}^{\mathrm{src}}+\sin^{2}\theta_{12}\,F_{\bar{\nu}_{X}}^{\mathrm{src}}&\text{(normal ordering)}\\ F_{\bar{\nu}_{X}}^{\mathrm{src}}&\text{(inverted ordering)}\end{cases}.$ (2)
@@
Different EOS models – LS180 ($M_{s=4}^{\rm max}=2.0,M_{\odot}$), LS220 ($M_{s=4}^{\rm max}=2.1,M_{\odot}$), DD2F RDF-1.7 ($M_{s=3}^{\rm max}=2.1,M_{\odot}$), SFHo ($M_{s=4}^{\rm max}=2.3,M_{\odot}$), Togashi ($M_{s=4}^{\rm max}=2.4,M_{\odot}$), Shen-TM1e ($M_{s=4}^{\rm max}=2.5,M_{\odot}$) and Shen-TM1 ($M_{s=4}^{\rm max}=2.6,M_{\odot}$) – are represented by symbols with varying shades and sizes. Most simulation results cluster around or below the 50% confidence boundary, indicating good consistency with observational constraints, particularly at lower neutrino energies. The Shen-TM1 EOS, associated with the largest maximum PNS mass ($2.6M_{\odot}$), predicts higher total neutrino energies, yet these results also fall within the observational limits. This clearly demonstrates how neutrino emission properties depend on the chosen EOS model.